-(4a-7)-5a=10+a
I have to solve for a. I've done TONS of problems like this before, but this one I just can't seem to figure out for some reason.

Answer by checkley71(8403)

-(4A-7)-5A=10+A
-4A+7-5A=10+A
-4A-5A-A=10-7
-10A=3
A=3/-10
A=-.3 ANSWER.
PROOF
-(4*-.3-7))-5*-.3=10-.3
-(-1.2-7)+1.5=9.7
-(-8.2)+1.5)=9.7
8.2+1.5=9.7
9.7=9.7